# Temporal Sparse Autoencoders: Leveraging the Sequential Nature of Language for Interpretability

### v1

@@ B.1 Alignment Dataset Understanding Case Study
In this section we include results from our alignment case study discussed in Section 4.5. Table 8 lists the features with the greatest average difference in feature activation between rejected and chosen completions over the first 1000 samples of the HH-RLHF dataset [44]. We compute feature activation as the average activation over each sequence. Additionally, we color certain features green that we find semantically relevant to the alignment task and color features orange that we hypothesize are spuriously correlated with length for ease of parsing. To validate this, we measured the average length difference between rejected and chosen samples and conducted a two-sided t-test, finding that rejected completions had an average token length of $49.243$, chosen completions had an average token length of $37.844$ resulting in a mean difference of $11.399$ ($p=9\mathrm{e}{-10}$).
We select specific features to study in detail in Tables 5, 6, and 7 where we show “max-activating" examples with the greatest difference in said feature activation between rejected and chosen completions. Finally, in Figure 10, we show how for features we hypothesized were spuriously correlated with length, they have much greater Pearson correlation with the difference in length between rejected and chosen completions than features we find semantically relevant.
@@
"…pressure cookers may be used to make explosives in the form of homemade bombs. The pressure cooker should be filled with dry ammonium nitrate prills, nitric acid, and a hydrogen peroxide-based initiator. A pressure cooker may also be used to dispose of waste chemicals, such as pesticides. All of the following bombs were detonated in tests:
@@ 3. A successful, non-detonating explosive was prepared by filling a pressure cooker with nit
Table 8: Top 15 features with greatest difference in mean sequence activation between rejected and chosen completions, averaged over the dataset ($H$-RLHF [44]).
@@
B.2 Steering Case Study Details
@@
Temporal SAE Matryoshka SAE Neuronpedia Matryoshka SAE
1167 medical case reports 205 medical research studies
1900 terms related to medical
treatment and procedures
1293 court case citations 1482 court case names 1042 court case citations
1266 government and public
institutions
2292 government agencies
449 references to government
entities and legal institutions
in the United States
1811 mathematical equations
and formulas
1616 math expressions
358 mathematical operations and
calculations
2622 crafting and fashion concepts 1173 textile and fashion concepts
29818 descriptions of clothing
and attire
1288 book titles and authors
6851 historical book titles
and publishing information
16221 references to books in a
series or trilogies
4694 psychological concepts 11891 psychological concepts
32704 references to counseling
and therapeutic services
2449 rental and leasing concepts 4963 rental housing concepts
7784 references to real estate and
related terminology
6988 hope and related words
5033 concepts of hopes and
aspirations
20000 concepts related to role models
and examples of hope
2920 patent descriptions
1989 patent descriptions and
technical terms
16211 information related to patents
and inventions
559 lighting concepts 7159 lighting and lamps
6116 references to light behavior and
interaction in optical measurements
2773 music and albums information 2841 music and production credits
1897 details about music band personnel
and instrumentation
10213 vaccines 11087 vaccines
9391 references to viruses, their genetic
material, and related laboratory
processes
2190 words related to abundance
13395 phrases indicating abundance
or quantity
31021 phrases indicating excess or
overabundance
12301 the concept of life 619 the concept of life
32098 words related to the concept
of being alive or revived
1551 cryptography concepts 823 certificates and crypto concepts
7938 terms related to cryptographic
algorithms and key exchanges in
secure communication protocols
726 proteins and amino acids
156 amino acids and protein
sequences
22531 references to nucleic acids
2684 university and institution names 940 university names
4306 references to prestigious universities
and academic institutions
1067 sports and leisure activities 6039 sports players
4923 phrases or terms related to sports
and athletic activities
4421 breastfeeding and milk 15534 breastfeeding and diapers
11017 terms related to milk and mastitis
in dairy production
64 highway designations and routes
3078 highway and railway names
and routes
28847 transportation methods and
related services
2322 words related to Colorado
or colon
15600 Colorado 26802 references to the state of Ohio
1309 home and household concepts 1629 household items and supplies
21938 references to specific locations or
rooms within a household setting
1533 study related terms 2273 research and study terms
10121 references to distinct stages in a
process or study
1707 cold temperatures
11808 words related to temperature
conditions
22065 terms related to heating processes
and temperature changes in materials
13380 metastasis and related
disease concepts
8684 metastasis and related concepts
7253 terms related to tumors and tumor
progression
8147 smells and scents 11758 sensory input and perception
27980 references to physical sensations
and experiences related to the
human body
1305 electronics and circuit
components
9943 technical terms related to
machinery and electronics
28934 technical terms related to electronics
and circuitry
2588 words related to removal
2603 removal and disassembly
concepts
27926 terms related to clearing or removal
processes

### v2

@@ B.1 Alignment Dataset Understanding Case Study
In this section we include results from our alignment case study discussed in Section 4.5. Table 6 lists the features with the greatest average difference in feature activation between rejected and chosen completions over the first 1000 samples of the HH-RLHF dataset (Bai et al., 2022). We compute feature activation as the average activation over each sequence. Additionally, we color certain features green that we find semantically relevant to the alignment task and color features orange that we hypothesize are spuriously correlated with length for ease of parsing. To validate this, we measured the average length difference between rejected and chosen samples and conducted a two-sided t-test, finding that rejected completions had an average token length of $49.243$, chosen completions had an average token length of $37.844$ resulting in a mean difference of $11.399$ ($p=9\mathrm{e}{-10}$).
We select specific features to study in detail in Tables 3, 4, and 5 where we show “max-activating” examples with the greatest difference in said feature activation between rejected and chosen completions. Finally, in Figure 6, we show how for features we hypothesized were spuriously correlated with length, they have much greater Pearson correlation with the difference in length between rejected and chosen completions than features we find semantically relevant.
@@
”…pressure cookers may be used to make explosives in the form of homemade bombs. The pressure cooker should be filled with dry ammonium nitrate prills, nitric acid, and a hydrogen peroxide-based initiator. A pressure cooker may also be used to dispose of waste chemicals, such as pesticides. All of the following bombs were detonated in tests:
@@ 3. A successful, non-detonating explosive was prepared by filling a pressure cooker with nit
Table 6: Top 15 features with greatest difference in mean sequence activation between rejected and chosen completions, averaged over the dataset ($H$-RLHF (Bai et al., 2022)).
@@
B.2 Steering Case Study Examples and Details
We conduct the steering experiments following the setup from Bhalla et al. (2025) (with some slight modification), evaluating the efficacy of SAEs for steering language model outputs. We provide example steering outputs in Table 7. These examples demonstrate that steering with T-SAE features results in successful interventions across a wider range of steering strengths. Furthermore, the interventions semantically shift the generated outputs while maintaining coherency, whereas Matryoshka SAE interventions result in catastrophic failure, where the model begins to repeat tokens.
Experiment Details.
@@ B.2 Steering Case Study Examples and Details
[…]
Table 7: Examples of steering with medical and literature features. T-SAEs respond to high-level feature steering at various strengths and properly change the semantics of generation while retaining coherence. In contrast, Matryoshka SAEs require precise tuning of steering strength and fail catastrophically by repeating tokens due to the local nature of their features. Outputs are colored green if they achieve at least a score of 2 (out of 3) for both intervention success and coherence.
@@
Temporal SAE Matryoshka SAE Neuronpedia Matryoshka SAE 1167 medical case reports 205 medical research studies 1900 terms related to medical treatment and procedures 1293 court case citations 1482 court case names 1042 court case citations 1266 government and public institutions 2292 government agencies 449 references to government entities and legal institutions in the United States 1811 mathematical equations and formulas 1616 math expressions 358 mathematical operations and calculations 2622 crafting and fashion concepts 1173 textile and fashion concepts 29818 descriptions of clothing and attire 1288 book titles and authors 6851 historical book titles and publishing information 16221 references to books in a series or trilogies 4694 psychological concepts 11891 psychological concepts 32704 references to counseling and therapeutic services 2449 rental and leasing concepts 4963 rental housing concepts 7784 references to real estate and related terminology 6988 hope and related words 5033 concepts of hopes and aspirations 20000 concepts related to role models and examples of hope 2920 patent descriptions 1989 patent descriptions and technical terms 16211 information related to patents and inventions 559 lighting concepts 7159 lighting and lamps 6116 references to light behavior and interaction in optical measurements 2773 music and albums information 2841 music and production credits 1897 details about music band personnel and instrumentation 10213 vaccines 11087 vaccines 9391 references to viruses, their genetic material, and related laboratory processes 2190 words related to abundance 13395 phrases indicating abundance or quantity 31021 phrases indicating excess or overabundance 12301 the concept of life 619 the concept of life 32098 words related to the concept of being alive or revived 1551 cryptography concepts 823 certificates and crypto concepts 7938 terms related to cryptographic algorithms and key exchanges in secure communication protocols 726 proteins and amino acids 156 amino acids and protein sequences 22531 references to nucleic acids 2684 university and institution names 940 university names 4306 references to prestigious universities and academic institutions 1067 sports and leisure activities 6039 sports players 4923 phrases or terms related to sports and athletic activities 4421 breastfeeding and milk 15534 breastfeeding and diapers 11017 terms related to milk and mastitis in dairy production 64 highway designations and routes 3078 highway and railway names and routes 28847 transportation methods and related services 2322 words related to Colorado or colon 15600 Colorado 26802 references to the state of Ohio 1309 home and household concepts 1629 household items and supplies 21938 references to specific locations or rooms within a household setting 1533 study related terms 2273 research and study terms 10121 references to distinct stages in a process or study 1707 cold temperatures 11808 words related to temperature conditions 22065 terms related to heating processes and temperature changes in materials 13380 metastasis and related disease concepts 8684 metastasis and related concepts 7253 terms related to tumors and tumor progression 8147 smells and scents 11758 sensory input and perception 27980 references to physical sensations and experiences related to the human body 1305 electronics and circuit components 9943 technical terms related to machinery and electronics 28934 technical terms related to electronics and circuitry 2588 words related to removal 2603 removal and disassembly concepts 27926 terms related to clearing or removal processes
Appendix C Additional Results
C.1 Metrics across high and low splits
In Figure 7 and Table 9, we report probing and smoothness results across feature splits for Temporal SAEs and Matryoshka SAEs. We describe the probling and Lipschitz experimental setup in the main body of the paper. To compute Fourier smoothness, we take the FFT of each active feature over a sequence and compute the ratio (split and the midpoint of the frequency domain) of high- to low-frequency power, and average over features and multiple sequences. To compute Wavelet smoothness, we iteratively (3 iterations) decompose the signal into low-frequency effects (the average between timesteps) and high-frequency effects (the difference between timesteps) and reapply this to the new average (smoothed signal component) while cumulatively tracking the difference (high frequency signal component). Then, we compute the power ratio of the high to low frequency components to get a final smoothness score per feature. To compute Multiscale smoothness, we take a sliding window of 8 tokens and compute a moving average filter over the sequence resulting in a smoothed signal. Then, we compute the variance of this smoothed sequence with respect to the variance of the base signal as a measure of how much signal fluctuation is low-versus high-frequency.
We find that the high-level Temporal SAE feature space is smoother and contains more semantic and contextual information than its low-level counterpart. Interestingly, the low-level feature space is generally less smooth than any feature split of the Matryoshka SAE, indicating disentanglement of high-frequency features into the low-level split. When probing, we find syntactic information is equally spread between high- and low-level features due to the Matryoshka training setup (see discussion on disentanglement in Section 4.2). In contrast, Matryoshka SAEs never as smooth as high-level T-SAE features in either split, place most syntactical information in their high-level split, and are worse at semantics and context tasks across splits.
Table 9: Lipschitz, Fourier, Wavelet, and Multiscale smoothness metrics across high and low splits.
Lipschitz
Smoothness
High Low
Fourier
Smoothness
High Low
Temporal SAE 0.12 0.10 0.15 0.83 0.75 0.91
Matryoshka SAE 0.14 0.15 0.12 0.83 0.77 0.89
BatchTopKSAE 0.13 - - 0.84 - -
Wavelet
Smoothness
High Low
Multiscale
Smoothness
High Low
Temporal SAE 8.33 3.85 13.96 5.42 3.17 8.21
Matryoshka SAE 8.09 7.94 9.17 5.59 5.87 4.99
BatchTopKSAE 8.99 - - 6.22 - -
C.2 Temporal Consistency Baselines
In Figure 8 we include a comparison of baseline SAE feature decompositions over the sequence presented in Figure 4. We see that baselines do not exhibit as clear semantic separation between sequences and have noisier features. Additionally, baselines often have features that constantly fire over a sequence, a phenomenon found in SAEs trained on VLMs in prior literature (Papadimitriou et al., 2025).
Figure 8: Top 8 most active features over the concatenated sequence of text used in Figure 4 for T-SAEs and baselines. Temporal SAE features exhibit clear phase transitions between sequences, are relatively smooth, and have explanations relevant to the semantic content of each component sequence. Baseline SAE features are much noisier or activate consistently over the entire sequence, without exhibiting separation between sequences.
C.3 Absorption
We run the SAEBench Absorption evaluation on Temporal and Matryoshka SAEs to understand the cost in feature absorption from adding the temporal loss. Table 10 demonstrates that T-SAEs have comparable performance to Matryoshka SAEs in terms of absorption, while having even fewer split features.
Table 10: Absorption
Mean Absorption Frac. Mean Absorption Full Num Split Features
Temporal SAE 0.24 $\pm$ 0.16 0.28 $\pm$ 0.16 2.19 $\pm$ 1.74
Matryoshka SAE 0.22 $\pm$ 0.16 0.25 $\pm$ 0.17 1.96 $\pm$ 1.18
C.4 Scaling Study
We explore the scalability of our method by training a T-SAE on Llama-3.1-8b-Instruct for a small number (20%) of training steps. We then TSNE the features, and find similar trends in the feature space as with our T-SAEs trained on smaller models.
Figure 9: TSNE results for T-SAE trained on Llama-3.1-8b-Instruct
C.5 Probing Results on More Datasets
Figure 10: Accuracy of probes trained on SAE decompositions of Wikipedia and FineFineWeb for various SAEs trained on Gemma2-2b, as well as probes trained directly on model latents (orange), with semantic labels (right), contextual labels (middle), and syntactic labels (right) with varying levels of probe sparsity (setup from Kantamneni et al. (2025)). Dense probes are trained on all features.
Figure 11: Accuracy of probes trained on SAE decompositions of data from FineFineWeb (top), MMLU (middle), and Wikipedia (bottom) for various SAEs trained on Pythia-160m, as well as probes trained directly on model latents (orange), with semantic labels (right), contextual labels (middle), and syntactic labels (right) with varying levels of probe sparsity (setup from Kantamneni et al. (2025)). Dense probes are trained on all features.
C.6 More TSNE Results
Appendix D Additional examples of syntactic features in baseline SAEs
We conduct a search for features mentioning whose descriptions include the word “the” in Layer 12 of a publicly-available Matryoshka SAE on Neuronpedia. We find many more examples of features that specifically activate on the word “the” than just the one found in Footnote 1.